OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 · x + 7 · 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (anonymous):

can someone help me

alones (alones):

sure it's B

OpenStudy (anonymous):

okay thank u

imqwerty (imqwerty):

1 Simplify 5+9 to 14 7×14=7x+7×9 2 Simplify 7×14 to 98 98=7x+7×9 3 Simplify 7×9 to 63 98=7x+63 4 Subtract 63 from both sides 98−63=7x 5 Simplify 98−63 to 35 35=7x 6 Divide both sides by 7 35/7=x 7 Simplify 35/7 to 5 5=x